A company expected its annual overhead costs to be $900,000 and direct labor costs to be $1,000,000. Actual overhead was $870,000, and actual labor costs totaled $1,100,000. How much is the company's predetermined overhead rate to the nearest cent?

$0.87

$0.82

$0.79

$0.90
Answer #1
Concepts and reason

Job order Costing: It is a method of cost accounting in which cost is collected and accumulated for each job, work order, or project separately. Especially the job order costing is followed in organizations where customized goods are produced.

Pre-Determined Overhead rate: The cost accountants estimate a rate for all the overhead costs to compute manufacturing overhead. That allocation rate, which is computed as a ratio of total estimated overhead costs to total estimated quantity, is known as predetermined overhead rate.

Fundamentals

Pre-Determined Overhead rate: The predetermined overhead allocation rate is used to allocate the total overhead costs to the direct labor costs.

Predeterminedoverheadrate}=TotalestimatedoverheadcostsDirectlaborcosts\left. \begin{array}{l}\\{\rm{Predetermined }}\\\\{\rm{overhead rate}}\\\end{array} \right\} = \frac{{{\rm{Total}}\,{\rm{estimated}}\,{\rm{overhead costs}}}}{{{\rm{Direct labor costs}}}}

Overhead costs: An overhead cost is an ongoing cost of operating a business. The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing overhead costs or indirect costs.

Direct labor costs: The labor cost includes an amount of wages or salaries paid to employees who physically produce goods or services. Labor refers to the actual work that the employees do to produce products. Labor costs refer to the amount of money that the employer pays the employees to complete the work.

Use the data to find out estimated overhead and direct labor costs.

Estimated overhead and direct labor costs directly in the problem as $900,000 and $1,000,000.

Calculate the Pre-Determined overhead rate as shown below:

Predeterminedoverheadrate}=TotalestimatedoverheadcostsDirectlaborcosts=$900,000$1,000,000=$0.90\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Predetermined }}\\\\{\rm{overhead rate}}\\\end{array} \right\} = \frac{{{\rm{Total}}\,{\rm{estimated}}\,{\rm{overhead costs}}}}{{{\rm{Direct labor costs}}}}\\\\ = \,\frac{{\$ 900,000}}{{\$ 1,000,000}}\\\\ = \,\$ 0.90\\\end{array}

Therefore, predetermined overhead rate is $0.90.

Ans:

Pre-determined overhead rate is nearest to $0.90.